OpenStudy (anikate):

Can someone plz help me find a fasetr way of finding out: (3x-y)^2-(5x-3y)^2 Plz try to tell me the fastest way.

OpenStudy (akshay_budhkar):

a^2-b^2= (a+b)(a-b) use that formula!

OpenStudy (anikate):

Yeah that's the problem how do I use that formula ?

OpenStudy (akshay_budhkar):

subsitute a=3x-y and b=5x-3y in the above formula

OpenStudy (anikate):

the a is squared so is the queston